OpenStudy (anonymous):

x^2+22x+? what would make this a perfect square trinomial

jigglypuff314 (jigglypuff314):

the "c" term in a perfect square trinomial is Half of the "b" squared c = (b/2)^2 in this case, b = 22

OpenStudy (anonymous):

so the ? is equal to 22?

jigglypuff314 (jigglypuff314):

no quite! :) b=22 plug that into (b/2)^2 = (22/2)^2 = "?"

OpenStudy (anonymous):

11^2??

jigglypuff314 (jigglypuff314):

almost there! :) 11^2 = 11*11 =

OpenStudy (anonymous):

121

jigglypuff314 (jigglypuff314):

there you go! tah dahhhhh